# Motion-aware Contrastive Learning for Temporal Panoptic Scene Graph Generation

### v2

@@ Introduction
The advent of autonomous agents, intelligent systems, and robots warrants a comprehensive understanding of real-world environments (Ma et al. 2022; Driess et al. 2023; Raychaudhuri et al. 2023; Cheng et al. 2022; Li et al. 2023b, a). Such understanding encompasses beyond merely recognizing individual entities, but also a sophisticated understanding of their relationships. To construct a detailed understanding, scene graph generation (SGG) research (Li, Yang, and Xu 2022; Sudhakaran et al. 2023; Nag et al. 2023; Wang et al. 2024a) has sought to provide relational perspective on scene understanding. In SGG frameworks, scene graphs utilize nodes to represent entities and edges to represent relationships, constructing a comprehensive and structured understanding of visual scenes.
@@
To encourage temporal representation learning, current research (Nguyen et al. 2023b; Liu et al. 2022; Zhou, Liu, and Wang 2023) uses contrastive learning for videos. However, they mainly seek to force two clips from the same video to be close together. As such, they mostly capture the semantics of visual scenes and disregard motions (Chen et al. 2020). Moreover, instead of paying attention to precise entity localization, they work upon frame-level representations. This would inadvertently inject motions from non-target entities into visual representations, which might not benefit relation classification in panoptic scene graph generation.
In this paper, to encourage representation learning to capture motion patterns for temporal panoptic scene graph generation, we propose a novel contrastive learning framework that focuses on mask tubes of the segmented entities. First, we force a mask tube and the one of similar subject-relation-object but of a different video to obtain close representations. Since positive mask tubes originate from distinct video clips, the model cannot rely upon visual semantics to optimize the contrastive objective, but instead depends on the motion trajectory evolution, which is our target component for representation learning. Second, we propel negative mask tubes generated by temporally shuffling the original tubes. Moreover, we also push apart representations of mask tubes from the same video but belonging to different triplets. Because mask tubes of different triplets from a common video with close visual features are separated from each other, we once again motivate the model to generate representations that are less reliant upon visual semantics but motion-sensitive features. In addition, the visually similar negative mask tubes can play a role as hard negative samples, thus accelerating the contrastive learning process (Chen, Zheng, and Song 2024).
@@
Video representation learning. Video representation learning has gained popularity in recent years (Nguyen et al. 2024d, f). Most approaches can be categorized into two groups: pretext-based and contrastive-based. Pretext-based methods mostly leverage pretext learning tasks such as optical flow prediction (Dong and Fu 2024; Davtyan, Sameni, and Favaro 2023) and temporal order prediction (Shen et al. 2024; Ren et al. 2024). However, these tasks are considerably influenced by low-level features and incapable of delving into high-level semantics of the video (Nguyen et al. 2024c). The contrastive-based methods primarily construct positive samples (Nguyen and Luu 2021; Wu et al. 2023; Nguyen et al. 2024a, 2022; Wu et al. 2024; Nguyen et al. 2024e, 2023a) by sampling video clips from the same video (Nguyen et al. 2024b; Liu et al. 2024), use various frame-based data augmentation techniques (Wang et al. 2024b; Song et al. 2024; Rosa 2024), etc., thereby increasing the similarity of positive pairs while simultaneously decreasing the one of negative pairs. Nevertheless, these methods focus on the whole video frames and are not suitable for object masks tracked across the temporal dimension.
@@
Given a video, our aim is to select mask tubes that carry strong motion for shuffling. To measure the motion of the mask tube, we utilize optical flow edges (Xiao, Tighe, and Modolo 2021). We estimate flow edges via employing a Sobel filter (Sobel et al. 2022) onto the flow magnitude map and take the median over the flow edge pixels of the entity masks. Then, we select mask tubes whose the maximum value across the optical flow surpasses a threshold $\gamma$.
@@
In order to create harder negative samples, we choose negative mask tubes from the same video with the anchor. We create a multi-nomial distribution, where triplets that share more subject, relation, or object categories with the anchor will be more likely to be drawn. Hence, our negative samples can hold close visual semantics with the anchor sample, and increase the likelihood that the model depends on motion semantics to push them apart. From contrastive learning perspective, these samples form hard negative samples to accelerate the learning process (Chen, Zheng, and Song 2024).
@@
$\begin{split}\text{s.t}\quad\Pi(\mathbf{a},\mathbf{b})=\{\mathbf{T}\in\mathbb{R}^{T_{i}\times T_{j}}_{+}\mid\mathbf{T}\mathbf{1}_{T_{i}}\leq\mathbf{a},\mathbf{T}^{\top}\mathbf{1}_{T_{j}}\leq\mathbf{b},\\ \mathbf{1}_{T_{i}}^{\top}\cdot\mathbf{T}\cdot\mathbf{1}_{T_{j}}=s,\quad 0\leq s\leq\min(T_{i},T_{j})\},\end{split}\vspace{-5pt}$ (8)
where $c$ denotes a pre-defined distance between two vectors. We implement the cost distance $c\left(\mathbf{h}_{i,k},\mathbf{h}_{j,l}\right)=1-\frac{\mathbf{h}_{i,k}\cdot\mathbf{h}_{j,l}}{||\mathbf{h}_{i,k}||_{2}||\mathbf{h}_{i,k}||_{2}}$ as the cosine distance. As the exact optimization over the transport plan $\mathbf{T}$ is intractable, we adopt the Sinkhorn-based algorithm to estimate $\mathbf{T}$. We delineate the algorithm to calculate the distance in Algorithm 1. To turn the distance into similarity value, we take its negative value and add to a pre-defined margin $\alpha$:
$\text{sim}\left(\mathbf{h}^{a},\mathbf{h}^{p}\right)=\alpha-d_{\text{OT}}.\vspace{-10pt}$ (9)
@@
Implementation details. For fair comparison, we experiment our contrastive framework with both IPS+T and VPS as segmentation module for panoptic video scene graph generation. In the former case, we leverage the UniTrack tracker (Wang et al. 2021) combined with Mask2Former model (Cheng et al. 2022), which is initialized from the best-performing COCO-pretrained weights and fine-tuned for 8 epochs using AdamW optimizer with a batch size of 32, learning rate of 0.0001, weight decay of 0.05, and gradient clipping with a max L2 norm of 0.01. In the latter case, we utilize Video K-Net (Li et al. 2022), also initialized from COCO-pretrained weights and fine-tuned with the same strategy as IPS+T. In the relation classification step, we conduct fine-tuning with a batch size of 32, employing the Adam optimizer with a learning rate of 0.001. For 4D panoptic scene graph generation, we adopt the PSG4DFormer baseline. To work with RGB-D and point cloud videos, we use an ImageNet pretrained on ResNet-101 (Russakovsky et al. 2015) and the DKNet (Wu et al. 2022) as the visual encoder, respectively. We fine-tune the segmentation module for RGB-D and point cloud videos for 12 and 200 epochs, respectively. We use additional 100 epochs to train the relation classification module. Based on validation, we adopt a threshold $\gamma=9.0$ and a margin $\alpha=10.0$. We set the maximum number of iterations $N_{\text{iter}}$ to 1,000.

### v3

@@ Introduction
The advent of autonomous agents, intelligent systems, and robots warrants a comprehensive understanding of real-world environments (Ma et al. 2022; Driess et al. 2023; Raychaudhuri et al. 2023; Cheng et al. 2022; Li et al. 2023b, a). Such understanding encompasses beyond merely recognizing individual entities, but also a sophisticated understanding of their relationships. To construct a detailed understanding, scene graph generation (SGG) research (Li et al. 2022b; Sudhakaran et al. 2023; Nag et al. 2023; Wang et al. 2024a) has sought to provide relational perspective on scene understanding. In SGG frameworks, scene graphs utilize nodes to represent entities and edges to represent relationships, constructing a comprehensive and structured understanding of visual scenes.
@@
To encourage temporal representation learning, current research (Nguyen et al. 2023b; Liu et al. 2022; Zhou et al. 2023) uses contrastive learning for videos. However, they mainly seek to force two clips from the same video to be close together. As such, they mostly capture the semantics of visual scenes and disregard motions (Chen et al. 2020). Moreover, instead of paying attention to precise entity localization, they work upon frame-level representations. This would inadvertently inject motions from non-target entities into visual representations, which might not benefit relation classification in panoptic scene graph generation.
In this paper, to encourage representation learning to capture motion patterns for temporal panoptic scene graph generation, we propose a novel contrastive learning framework that focuses on mask tubes of the segmented entities. First, we force a mask tube and the one of similar subject-relation-object but of a different video to obtain close representations. Since positive mask tubes originate from distinct video clips, the model cannot rely upon visual semantics to optimize the contrastive objective, but instead depends on the motion trajectory evolution, which is our target component for representation learning. Second, we propel negative mask tubes generated by temporally shuffling the original tubes. Moreover, we also push apart representations of mask tubes from the same video but belonging to different triplets. Because mask tubes of different triplets from a common video with close visual features are separated from each other, we once again motivate the model to generate representations that are less reliant upon visual semantics but motion-sensitive features. In addition, the visually similar negative mask tubes can play a role as hard negative samples, thus accelerating the contrastive learning process (Chen et al. 2024).
@@
Video representation learning. Video representation learning has gained popularity in recent years (Nguyen et al. 2024e, d). Most approaches can be categorized into two groups: pretext-based and contrastive-based. Pretext-based methods mostly leverage pretext learning tasks such as optical flow prediction (Dong and Fu 2024; Davtyan et al. 2023) and temporal order prediction (Shen et al. 2024; Ren et al. 2024). However, these tasks are considerably influenced by low-level features and incapable of delving into high-level semantics of the video (Nguyen et al. 2024c). The contrastive-based methods primarily construct positive samples (Nguyen and Luu 2021; Wu et al. 2023; Nguyen et al. 2024a, 2022; Wu et al. 2024; Nguyen et al. 2024f, 2023a) by sampling video clips from the same video (Nguyen et al. 2024b; Liu et al. 2024), use various frame-based data augmentation techniques (Wang et al. 2024b; Song et al. 2024; Rosa 2024), etc., thereby increasing the similarity of positive pairs while simultaneously decreasing the one of negative pairs. Nevertheless, these methods focus on the whole video frames and are not suitable for object masks tracked across the temporal dimension.
@@
Given a video, our aim is to select mask tubes that carry strong motion for shuffling. To measure the motion of the mask tube, we utilize optical flow edges (Xiao et al. 2021). We estimate flow edges via employing a Sobel filter (Sobel et al. 2022) onto the flow magnitude map and take the median over the flow edge pixels of the entity masks. Then, we select mask tubes whose the maximum value across the optical flow surpasses a threshold $\gamma$.
@@
In order to create harder negative samples, we choose negative mask tubes from the same video with the anchor. We create a multi-nomial distribution, where triplets that share more subject, relation, or object categories with the anchor will be more likely to be drawn. Hence, our negative samples can hold close visual semantics with the anchor sample, and increase the likelihood that the model depends on motion semantics to push them apart. From contrastive learning perspective, these samples form hard negative samples to accelerate the learning process (Chen et al. 2024).
@@
$\begin{split}\text{s.t}\quad\Pi(\mathbf{a},\mathbf{b})=\{\mathbf{T}\in\mathbb{R}^{T_{i}\times T_{j}}_{+}\mid\mathbf{T}\mathbf{1}_{T_{i}}\leq\mathbf{a},\mathbf{T}^{\top}\mathbf{1}_{T_{j}}\leq\mathbf{b},\\ \mathbf{1}_{T_{i}}^{\top}\cdot\mathbf{T}\cdot\mathbf{1}_{T_{j}}=s,\quad 0\leq s\leq\min(T_{i},T_{j})\},\end{split}\vskip-5.0pt$ (8)
[…]
$\text{sim}\left(\mathbf{h}^{a},\mathbf{h}^{p}\right)=\alpha-d_{\text{OT}}.\vskip-10.0pt$ (9)
@@
Implementation details. For fair comparison, we experiment our contrastive framework with both IPS+T and VPS as segmentation module for panoptic video scene graph generation. In the former case, we leverage the UniTrack tracker (Wang et al. 2021) combined with Mask2Former model (Cheng et al. 2022), which is initialized from the best-performing COCO-pretrained weights and fine-tuned for 8 epochs using AdamW optimizer with a batch size of 32, learning rate of 0.0001, weight decay of 0.05, and gradient clipping with a max L2 norm of 0.01. In the latter case, we utilize Video K-Net (Li et al. 2022a), also initialized from COCO-pretrained weights and fine-tuned with the same strategy as IPS+T. In the relation classification step, we conduct fine-tuning with a batch size of 32, employing the Adam optimizer with a learning rate of 0.001. For 4D panoptic scene graph generation, we adopt the PSG4DFormer baseline. To work with RGB-D and point cloud videos, we use an ImageNet pretrained on ResNet-101 (Russakovsky et al. 2015) and the DKNet (Wu et al. 2022) as the visual encoder, respectively. We fine-tune the segmentation module for RGB-D and point cloud videos for 12 and 200 epochs, respectively. We use additional 100 epochs to train the relation classification module. Based on validation, we adopt a threshold $\gamma=9.0$ and a margin $\alpha=10.0$. We set the maximum number of iterations $N_{\text{iter}}$ to 1,000.